# MSDiagnosis: An EMR-based Dataset for Clinical Multi-Step Diagnosis

### v1

@@ 5.2 Forward Inference
Specifically, given the admission record $E$ and question $Q_{o}$, we select ICL examples with similar semantics to $E$ through the following two steps, where $o$ is the total number of questions. First, we utilize the BGE Xiao et al. (2023) model to obtain the representations of $E$ and each sample $Y_{i}\ (1\leq i\leq u)$ in the training set, where $u$ is the total number of training set. Second, we calculate their cosine similarity and select the top $K$ samples with the highest similarity as ICL examples. In addition to using $E$ and the top $K$ samples as input to the LLM, we introduce a role definition and a predefined rule. An illustrative example of this process is presented in Fig. 5 in Appendix I. The role definition is phrased as “You are a professional doctor, and you need to complete a task related to diagnosis”. The rules for constraining the output format of the LLM refer to “The diagnostic results should be in the following format, so they can be directly loaded using the JSON.load() function”. Finally, the LLM would generate the answer $a_{o}$ corresponding to question $Q_{o}$.
5.3 Backward Inference and Reflection
After obtaining the forward inference results, this part aims to conduct backward inference, reflection, and refine the patient’s diagnosis. In this paper, we design specific rules to achieve this goal.
Specifically, given the forward inference diagnostic results $a_{o}$, we first perform backward inference from the diagnosis to the diagnostic criteria. In this step, we define backward inference rules that aim to guide the LLM to generate output that complies with pre-defined content and format constraints, as shown in Fig. 7 in Appendix I. For content constraints, we have “For each diagnosis in the diagnostic result, recall the representative medical history, symptoms, physical signs, and auxiliary examination results for the disease” (Rule 1). For format constraints, we consider “The recalled content should be in the following format: Medical History: <Recall the representative medical history for the disease; delete this item if not applicable>…” (Rule 2). Then, based on the results of the backward inference, we design reflection rules to review the diagnostic results, as shown in Fig. 8 in Appendix I. Finally, we combine the aforementioned backward inference, reflection results, and their criteria to let the model optimize the diagnostic results, with specific prompts shown in Fig. 6 in Appendix I.
F1 Macro-Recall Rouge-L Blue-1
Model Settings Pre DD Fin Pre Fin Pre Fin Pre Fin
@@ 5.3 Backward Inference and Reflection
MMed-Llama-3-8B 1-Shot 22.03 8.87 18.19 40.73 32.68 22.29 14.77 12.83 6.02
PULSE-20bv5 1-Shot 26.40 9.55 13.18 27.57 24.37 45.93 42.20 36.11 32.44
Llama3-OpenBioLLM-8B 1-Shot 00.41 9.90 0.14 18.62 10.19 11.56 7.30 8.57 5.50
@@
1-Shot 32.72 5.54 19.99 37.88 22.11 51.91 32.10 40.80 27.86
Llama-3.1-8B LoRA 20.59 7.86 16.32 17.04 14.06 25.07 19.76 18.69 15.00
1-Shot 34.71 9.97 30.92 41.79 40.79 53.68 55.44 47.72 46.73
@@
Other Methods
Human - 95.11 84.31 96.31 97.08 98.16 98.64 99.44 98.27 98.41
Table 3: Method comparisons on ReaDiagnosis (%). “Pre” refers to the primary diagnosis. “Fin” stands for the final diagnosis. “DD” means differential diagnosis. All results are the average of three runs. The second results are highlighted by underline.
@@ 6 Experiments
In this section, we first conduct extensive experiments to evaluate the effectiveness of our proposed framework. Next, we provide a detailed analysis to offer deeper insights into our framework.
@@ 6.1.2 Evaluation Details
For different types of problems, we introduce various evaluation metrics. Specifically, for primary diagnosis, differential diagnosis, and final diagnosis, we introduce entity $F1$ Liu et al. (2022) as the evaluation metric. For primary and final diagnosis criteria, we employ two types of evaluation metrics. First, we adapt $Rouge-L$ Lin (2004) and $BLEU$-1 Papineni et al. (2002) to measure the similarity between the generated text and the reference text. Second, we introduce the $Macro-Recall$ metric, which is calculated based on the key information in the answers. More details are shown in Appendix B.
@@ 6.1.3 Implementation Details
For all the open-source models mentioned above, we use their default hyperparameters. The further implementation details are listed in Appendix C.
@@ 6.2 Main Results
To verify the effectiveness of our proposed method, we compare it with all baselines on the ReaDiagnosis test set. The results are reported in Table 3.
From the table, we conclude that: 1) All LLMs perform poorly on ReaDiagnosis, with significant room for improvement compared to the human final $F1$ of 95.11%. For diagnostic-type questions, the best-performing model, GLM with LoRA, achieves a primary $F1$ score of 38.78% and a final $F1$ score of 35.00%. For diagnostic criteria questions, the best-performing model, Baichuan2 with LoRA, reaches a primary $Macro-Recall$ of 51.15% and a final $Macro-Recall$ of 52.58%. 2) Our method outperforms all baselines in the $Macro-Recall$, $Rouge-L$, and $Blue$-1 metrics, demonstrating the effectiveness of the proposed approach. 3) Among open-source General LLMs, the GLM and Baichuan models, after instruction fine-tuning, outperform those without instruction fine-tuning in diagnostic reasoning. Specifically, after instruction fine-tuning, the GLM shows improvements of 3.74% and 2.4% in $Pre-F1$ and $Fin-F1$ metrics, respectively. However, the performance of the LLaMA model declines after instruction fine-tuning, with $Pre-F1$ and $Fin-F1$ metrics decreasing by 12.13% and 3.67%, respectively. This decline may be due to the relatively small proportion of Chinese in LLaMA’s training corpus, which weakens the model’s Chinese language capabilities, and direct instruction fine-tuning may further reduce its performance.
@@ 6.3 Detailed Analysis
Method F1 Macro-Recall Rouge-L Blue-1
Pre DD Fin Pre Fin Pre Fin Pre Fin
w/o Backward inference 34.59 10.96 32.04 41.32 41.01 55.48 55.48 48.93 47.32
w/o Reflection 34.68 10.04 32.19 41.08 41.60 55.53 55.89 49.16 47.87
w/o Refinement 34.63 10.81 32.92 41.95 42.15 55.63 55.39 49.06 47.72
Ours 34.78 11.13 34.32 42.67 42.28 55.95 55.97 49.15 47.94
Table 4: The ablation results of our framework (%).
@@ 6.3.1 Ablation Study
In this section, we design two ablation experiments to analyze the multi-stage diagnostic process of ReaDiagnosis and our framework respectively. Specifically, for the ablation of the multi-stage diagnostic process, we sequentially remove the primary diagnosis, differential diagnosis, and both to evaluate the effectiveness of the final diagnosis. It is important to note that in this experiment, we employ a method where the LLM performs direct reasoning. The experimental results are shown in Table 5. For the ablation of our framework, we sequentially remove backward inference, reflection, and refinement, and assess the effectiveness of the remaining combinations. The results are presented in Table 4.
From the Table 5, we conclude that: 1) The primary diagnosis significantly influences the performance of the final diagnosis. This could be due to errors in the primary diagnosis propagating through the multi-step diagnostic process, leading to a reduction in the effectiveness of the final diagnosis. 2) The multi-step diagnostic process helps improve the interpretability of the final diagnosis. Specifically, by introducing a multi-step diagnosis, the final diagnosis is based on more criteria, thereby enhancing its interpretability. From the Table 4, we have the following observations: The removal of components backward inference, reflection, and refinement leads to a decline in performance, illustrating the effectiveness of these components. It is worth noting that when backward inference is removed, both the final $F1$ and final $Macro-Recall$ scores decrease significantly, by 1.28% and 1.27%, respectively. This indicates that backward inference effectively enhances the accuracy and interpretability of the final diagnosis.
F1 Macro-Recall Rouge-L Blue
Method Pre DD Fin Pre Fin Pre Fin Pre Fin
Primary diagnosis matches final diagnosis
w/o primary diagnosis - 8.05 44.46 - 42.66 - 59.53 - 52.30
w/o differential diagnosis 40.8 - 41.28 35.96 42.10 57.14 58.67 48.08 50.62
multi-step 38.11 8.23 39.33 35.94 42.96 55.94 58.62 46.88 49.21
Primary diagnosis differs from final diagnosis
w/o primary diagnosis - 9.80 21.17 - 29.72 - 49.66 - 42.35
w/o differential diagnosis 25.19 - 18.78 32.70 30.08 50.28 49.04 43.69 41.91
multi-step 23.99 11.49 18.12 32.50 30.91 50.74 48.76 42.90 41.88
Table 5: The ablation results of the multi-step diagnostic process (%).
@@ 6.3.2 Case Study
To further investigate the limitations of our method, we analyze 100 error samples generated in ReaDiagnosis. After manual classification, the errors are categorized into four types: (a) lack of domain knowledge (42%), (b) Symptoms or examination results are confused with the disease. (30%), (c) diagnostic basis inconsistent with facts (18%), and (d) other (10%). To illustrate these error types more intuitively, we provide examples as shown in Fig. 4 in appendix H.
@@ 7 Related Work
The related work in this paper can be categorized into two groups: clinical diagnosis on EMRs and prompting strategies of LLM.
@@ 7.1 Clinical Diagnosis on EMRs
Clinic diagnosis based on EMRs is crucial for improving healthcare quality and patient outcomes. However, existing datasets for EMR-based diagnosis primarily focus on single-step diagnosis, with diagnostic types generally divided into two categories: primary diagnosis and differential diagnosis. Primary diagnosis involves determining the likely disease based on the patient’s history and symptoms, or in combination with examination results. Most current research focuses on extracting abnormal features from medical records or integrating external medical knowledge to make a diagnosis Xu et al. (2024) Jia et al. (2024) Zhu et al. (2024). Differential diagnosis is a list of potential diseases that could cause the patient’s symptoms Adler-Milstein et al. (2021). It enables a more comprehensive evaluation of clinical EMRs, allowing for the identification of less obvious but critical conditions. Current studies often employ deep learning methods to extract features from medical records for differential diagnosis Zhou et al. (2024) Wu et al. (2023). In summary, previous research on EMR-based diagnosis has primarily focused on single-step diagnostic processes. However, this approach does not align with actual clinical diagnostic workflows.
@@ 7.2 Prompting Strategies of LLM
With the development of LLMs, many researchers have applied these models to medical downstream tasks. In these tasks, specific problems are typically addressed by designing prompts. These methods can be categorized into three types: Input-Output(IO) prompting, CoT prompting, and the DAC paradigm. IO prompting Yao et al. (2024) is a standard prompting strategy where input is combined with instructions and/or a few in-context learning(ICL) examples to generate a response. CoT prompting Wei et al. (2022) aims to emulate the step-by-step thought process humans use to tackle complex tasks, such as combinatorial reasoning and mathematical calculations. There are also various CoT variants, such as CoT with self-consistency (CoT-SC) prompting Wang et al. (2022), designed to address the limitations of CoT in exploration. The DAC paradigm Zhang et al. (2024) mainly refers to simply breaking down the input sequence into multiple sub-inputs to enhance LLM performance on certain specific tasks. Although these methods have shown promising results in some reasoning tasks, this reasoning approach is challenging to directly apply to medical diagnostic reasoning tasks.
8 Conclusion and Limitations
We propose ReaDiagnosis, a multi-step diagnostic reasoning dataset collected and annotated from medical websites. ReaDiagnosis addresses the weaknesses of existing datasets by constructing multi-stage diagnostic questions that better align with actual clinical diagnostic scenarios. On this dataset, we propose a simple and effective framework. We implement a series of closed-source and open-source LLMs and conduct extensive experimental analyses. The results demonstrate that ReaDiagnosis is a challenging dataset worthy of further exploration. Moreover, the method proposed in this paper shows effectiveness on this dataset. However, the dataset has a limitation: due to limited data sources, our medical records dataset exhibits an uneven distribution across different departments. This issue can be addressed through machine learning methods Cao et al. (2019) and data sampling strategies Chawla et al. (2002).
@@ 8 Conclusion and Limitations
Our data source adheres to ethical guidelines and respects copyright laws. The entire data collection process is free from copyright and privacy concerns. In addition, we have meticulously reviewed our dataset to ensure it does not contain any harmful content, including gender bias, racial discrimination, or inappropriate material.
@@ Appendix A Details of the Baseline
In the open-source Medical LLMs, we employ MMedLM Qiu et al. (2024), PULSE Xiaofan Zhang (2023), and Lmama3-OpenBioLLM Ankit Pal (2024) for comparison. Based on these models, we manually construct an example to serve as ICL. In the open-source General LLMs, we use Llama Touvron et al. (2023), glm GLM et al. (2024), and Baichuan Yang et al. (2023) for comparison. Based on these models, we design two settings: 1-shot reasoning and instruction tuning. In the first setting, we utilize the previously constructed example as ICL. In the second setting, we use parameter-efficient fine-tuning with LoRA Hu et al. (2021). The instruction data is generated by transforming the input and output from the training data. In the closed-source LLMs, we employ two settings: 1-shot and prompting methods, In the first approach, we utilize the same example as previously used. In the second setting, we consider comparing method COT Wei et al. (2022) and method DAC Zhang et al. (2024). COT Wei et al. (2022) uses “Let’s think step by step” to enhance the model’s reasoning ability for task-solving. DAC Zhang et al. (2024) adopts a simple divide-and-conquer prompting strategy, where the input sequence is simply divided into multiple sub-inputs, which can enhance the reasoning performance of the LLM. In other methods, We mainly consider manual answering methods. Specifically, we randomly select 100 samples from ReaDiagnosis and invite one college student (different from the annotation team in Section 3.2.2) to answer the questions.

### v3

@@ 5.2 Forward Inference
Specifically, given the admission record $E$ and question $Q_{o}$, where $o$ is the total number of questions, we select ICL examples with similar semantics to $E$ through the following steps. First, we utilize the BGE Xiao et al. (2023) model to obtain the representations of $E$ and each sample $Y_{i}\ (1\leq i\leq u)$ in the training set, where $u$ is the total number of the training set. Second, we calculate their cosine similarity and select the top $K$ samples with the highest similarity as ICL examples. In addition to using $E$ and the top $K$ samples as input to the LLM, we introduce a role definition and a predefined rule. The role definition is phrased as “You are a professional doctor, and you need to complete diagnosis task”. The rules for constraining the output format of the LLM refer to “The output format of the diagnostic results can be loaded directly using the JSON.load() function”. Finally, the LLM would generate the answer $a_{o}$ corresponding to question $Q_{o}$. The detailed prompt is shown in Appendix I.
[…]
F1 Macro-Recall Rouge-L BLEU-1 BERTScore
Model Settings Pre DD Fin Pre Fin Pre Fin Pre Fin Pre Fin
@@
MMed-Llama-3-8B 1-Shot 22.03 8.87 18.19 40.73 32.68 22.29 14.77 12.83 6.02 60.90 60.13
PULSE-20bv5 1-Shot 26.40 9.55 13.18 27.57 24.37 45.93 42.20 36.11 32.44 80.46 78.46
Apollo2-7B 1-Shot 33.61 8.47 26.78 37.01 33.15 51.25 44.91 37.00 33.10 80.70 78.50
@@
1-Shot 32.72 5.54 19.99 37.88 22.11 51.91 32.10 34.42 27.86 82.10 77.13
Llama-3.1-8B LoRA 20.59 7.86 16.32 17.04 14.06 25.07 19.76 18.69 15.00 70.97 70.20
@@
CoT 30.32 3.63 22.29 26.51 19.10 38.65 34.18 29.47 25.91 72.36 70.89
DAC 24.74 8.07 25.05 28.82 35.34 40.40 49.64 29.88 42.37 81.17 82.35
ChatGPT3.5-turbo Ours 34.60 8.23 23.29 39.38 39.02 51.54 47.19 47.04 45.05 82.27 83.31
1-Shot 31.00 10.10 28.92 34.61 35.82 53.28 53.61 44.91 45.45 84.39 84.56
CoT 33.79 7.06 31.28 27.63 26.31 48.95 48.67 38.66 37.55 78.63 79.03
DAC 26.95 10.17 24.90 23.18 24.26 52.33 51.96 40.42 39.86 80.83 81.13
[…]
Other Method
Human - 94.11 84.31 95.31 96.08 97.16 97.64 97.44 97.27 97.41 98.03 98.19
5.3 Backward Inference and Reflection
After obtaining the forward inference results, this part aims to conduct backward inference, reflection, and refine the patient’s diagnosis. In this section, we define some rules to achieve this goal.
Specifically, given the forward inference diagnostic results $a_{o}$, we first perform backward inference from the diagnosis to diagnostic criteria. In this step, we define backward inference rules to guide the LLM in generating outputs that comply with pre-defined content and format constraints. For content constraints, we have “For each diagnosis, recall the representative medical history, symptoms, physical signs, and examination results” (Rule 1). For format constraints, we consider “The recalled content should follow the format: Medical History: Recall the representative medical history for the disease; delete this item if not applicable” (Rule 2). Then, we design reflection rules to review the diagnoses. The specific reflection rule is: “If a diagnosis’s characteristics don’t align with the medical record, delete or revise it, and provide the rationale”. Finally, we combine the aforementioned backward inference, reflection results, and their criteria to let the model optimize the diagnostic results, with specific prompts shown in Appendix I.
@@ 6 Experiments
In this section, we conduct a series of experiments to evaluate the performance of LLMs on the MSDiagnosis and the effectiveness of our proposed framework. Additionally, we provide an analysis of the multi-step diagnostic process and the proposed framework.
@@ 6.1.2 Evaluation Details
For different types of problems, we introduce various evaluation metrics. Specifically, for primary diagnosis, differential diagnosis, and final diagnosis, we introduce entity $F1$ Liu et al. (2022) as the evaluation metric. For primary and final diagnosis criteria, we employ two types of evaluation metrics. First, we adapt $Rouge-L$ Lin (2004), $BERTScore$ Zhang* et al. (2020), and $BLEU$-1 Papineni et al. (2002) to measure the similarity between the generated text and the reference text. Second, we introduce the $Macro-Recall$ metric, which is calculated based on the key points in the answers. More details are shown in Appendix B.
@@ 6.1.3 Implementation Details
For all the open-source models mentioned above, we use their default hyperparameters. All experiments are conducted three times, and the average performance across three runs is computed. Further implementation details are listed in Appendix C.
Table 4: The ablation results of the multi-step diagnostic (%). w/o represents deleting the corresponding process.
F1 Macro-Recall Rouge-L BLEU-1 BERTScore
Method Pre DD Fin Pre Fin Pre Fin Pre Fin Pre Fin
Primary diagnosis matches final diagnosis
w/o primary diagnosis - 8.05 44.46 - 42.66 - 59.53 - 52.30 - 86.40
w/o differential diagnosis 40.8 - 41.28 35.96 42.10 57.14 58.67 48.08 50.62 85.55 86.08
multi-step 38.11 8.23 39.33 35.94 42.96 55.94 58.62 46.88 49.21 85.14 85.77
Primary diagnosis differs from final diagnosis
w/o primary diagnosis - 9.80 21.17 - 29.72 - 49.66 - 42.35 - 83.05
w/o differential diagnosis 25.19 - 18.78 32.70 30.08 50.28 49.04 43.69 41.91 83.62 83.03
multi-step 23.99 11.49 18.12 32.50 30.91 50.74 48.76 42.90 41.88 83.46 83.09
Table 5: The ablation results of our framework (%). w/o indicates the removal of the corresponding module.
Method F1 Macro-Recall Rouge-L BLEU-1 BERTScore
Pre DD Fin Pre Fin Pre Fin Pre Fin Pre Fin
w/o Backward inference 34.59 10.96 32.04 41.32 41.01 55.48 55.48 48.93 47.32 85.83 84.76
w/o Reflection 34.68 10.04 32.19 41.08 41.60 55.53 55.89 49.16 47.87 85.93 85.93
w/o Refinement 34.63 10.81 32.92 41.95 42.15 55.63 55.39 49.06 47.72 85.16 84.83
Ours 34.78 11.13 34.32 42.67 43.28 55.95 55.97 49.15 47.94 86.13 86.06
@@ 6.2 Main Results
To verify the effectiveness of our proposed method, we compare it with all baselines on the MSDiagnosis test set. The results are reported in Table 3.
From the table, we conclude that: 1) All LLMs perform poorly on MSDiagnosis, with a significant gap from the human final diagnosis $F1$ of 95.31%. Specifically, for the final diagnosis question, the best-performing model, glm4 with LoRA, achieves a final $F1$ of 38.78%. 2) Our framework outperforms all baselines in the final diagnosis $F1$ and $Blue$-1 metrics, demonstrating the effectiveness of the proposed framework. In particular, our framework improves the final $F1$ by 0.32% over glm4 with LoRA. 3) Among open-source general LLMs, the glm4 and Baichuan2 models outperform those without instruction fine-tuning in diagnostic reasoning after fine-tuning. Specifically, after fine-tuning, glm4 improves by 4.07% and 2.24% in primary $F1$ and final $F1$ metrics, respectively. However, the LlaMA’s performance declines after fine-tuning, with primary $F1$ and final $F1$ metrics decreasing by 12.13% and 3.67%, respectively. This decline likely results from the relatively small proportion of Chinese in LlaMA’s training corpus. Using Chinese data for instruction fine-tuning could further degrade its performance. 4) The model performs worse in differential diagnosis than in primary and final diagnoses. This is due to the increased complexity of the differential diagnosis task, which requires distinguishing between similar diseases and demands more specialized medical knowledge.
@@ 6.3 Detailed Analysis
Figure 4: Case study. The green (red) highlight indicates correct (incorrect) results. Purple marks lack of domain knowledge errors, blue marks symptom-disease confusion errors, and yellow marks diagnostic criteria inconsistent with facts.
@@ 6.3.1 Ablation Study
In this section, we design two ablation experiments to analyze the multi-step diagnostic process of MSDiagnosis and our framework respectively. Specifically, for the ablation of the multi-step diagnostic process, we sequentially remove the primary diagnosis, differential diagnosis, and both to evaluate the effectiveness of the final diagnosis. Notably, the LLM is used for direct reasoning in this experiment. The results are shown in Table 4. For the ablation of our framework, we sequentially remove backward inference, reflection, and refinement, assessing the effectiveness of the remaining components. The results are presented in Table 5.
From the Table 4, we conclude that: 1) The primary diagnosis significantly influences the final diagnosis’s performance, likely due to errors in the primary diagnosis propagating through the multi-step process, reducing the final diagnosis’s effectiveness. 2) The multi-step diagnostic process helps improve the interpretability of the final diagnosis. Specifically, by introducing a multi-step diagnosis, the final diagnosis is based on more criteria, thereby enhancing its interpretability. From the Table 5, it can be observed that when backward reasoning is removed, the $F1$ and $Macro-Recall$ scores for the final diagnosis decrease significantly, specifically by 2.28% and 2.27%, respectively. These findings suggest that backward reasoning effectively enhances both the accuracy and interpretability of the final diagnosis.
@@ 6.3.2 Case Study
To further investigate the limitations of our method, we analyze 100 error samples. After manual classification, the errors are categorized into three types: (a) lack of domain knowledge (42%), (b) Symptoms or examination results are confused with the disease. (30%), (c) diagnostic criteria inconsistent with facts (18%), along with other errors (10%). To illustrate these error types more intuitively, we provide examples as shown in Fig. 4.
@@ 7.1 Clinical Diagnosis on EMRs
Clinic diagnosis based on EMRs is crucial for improving healthcare quality and patient outcomes. However, existing datasets for EMR-based diagnosis primarily focus on single-step diagnosis, with diagnostic types generally divided into two categories: primary diagnosis and differential diagnosis. Primary diagnosis involves determining the likely disease based on the patient’s history and symptoms, or in combination with examination results. Most current research focuses on extracting abnormal features from medical records or integrating external medical knowledge to make a diagnosis Xu et al. (2024); Jia et al. (2024); Zhu et al. (2024). Differential diagnosis is a list of potential diseases that could cause the patient’s symptoms Adler-Milstein et al. (2021). It enables a more comprehensive evaluation of clinical EMRs, allowing for the identification of less obvious but critical conditions. Current studies often employ deep learning methods to extract features from medical records for differential diagnosis Zhou et al. (2024); Wu et al. (2023). In summary, previous research on EMR-based diagnosis has primarily focused on single-step diagnostic processes. However, this approach does not align with actual clinical diagnostic workflows.
@@ 7.2 Prompting Strategies of LLM
With the development of LLMs, many researchers have applied these models to medical tasks. These methods can be categorized into three types: IO prompting, CoT prompting, and the DAC paradigm. IO prompting Yao et al. (2024) is a standard prompting strategy where input is combined with instructions and/or a few ICLs to generate a response. CoT prompting Wei et al. (2022) aims to emulate the step-by-step thought process humans use to tackle complex tasks, such as combinatorial reasoning and mathematical calculations. There are also various CoT variants, such as CoT with self-consistency (CoT-SC) prompting Wang et al. (2022), designed to address the limitations of CoT in exploration. The DAC paradigm Zhang et al. (2024) mainly refers to simply breaking down the input sequence into multiple sub-inputs to enhance LLM performance on certain specific tasks. While these methods show promise in reasoning tasks, they are challenging to apply directly to medical diagnostic reasoning.
8 Conclusion
This paper introduces MSDiagnosis, a multi-step clinical diagnostic benchmark that is collected and annotated from medical websites. MSDiagnosis addresses the limitations of existing datasets by constructing multi-step diagnostic tasks that better align with actual clinical diagnostic scenarios. For this benchmark, we propose a simple and effective framework. We implement both closed-source and open-source LLMs and conduct extensive experiments. The results show that tasks in our benchmark effectively measure the multi-step clinical diagnostic abilities, and the framework proposed in this paper shows effectiveness on this benchmark.
Limitations
The limitations of our work are as follows. Firstly, due to limited data sources, our medical records dataset exhibits an uneven distribution across different departments. This issue can be addressed through machine learning methods Cao et al. (2019) and data sampling strategies Chawla et al. (2002). Secondly, the proposed framework requires multiple uses of the LLM, resulting in a longer inference time for multi-step diagnostic reasoning compared to direct reasoning.
@@ 8 Conclusion
The raw data used in this study comes from an open-source medical platform, which includes a large volume of EMRs. The platform permits data for research and education, as confirmed by prior studies Yim et al. (2024a); Li et al. (2024); Yim et al. (2024b). Throughout the data collection process, we strictly ensured the avoidance of copyright and privacy issues. Furthermore, we conducted a thorough review of the dataset to ensure it contains no harmful content, such as gender bias, racial discrimination, or other inappropriate materials.
@@ Appendix A Details of the Baseline
In this paper, we mainly describe several types of baseline methods, including open-source medical LLMs, open-source general LLMs, closed-source LLMs, and other method.
In the open-source medical LLMs, we employ MMedLM Qiu et al. (2024), PULSE Xiaofan Zhang (2023), and Apollo2-7B Zheng et al. (2024) for comparison. Based on these models, we manually construct an example to serve as ICL. The complete example is shown in Section J. We previously tested several medical models (HuatuoGPT2-7B Chen et al. (2024), DoctorGLM Xiong et al. (2023), and BianQue-2 Chen et al. (2023)), but they were excluded due to poor performance, as the length of our dataset approached their context window limits. According to our statistics, in the diagnostic tasks of this study, the average number of context tokens for the primary diagnosis is 3245.14, for the differential diagnosis is 3336.75, and for the final diagnosis is 4807.20.
In the open-source general LLMs, we use Llama-3.1 Touvron et al. (2023), glm4 GLM et al. (2024), and Baichuan2 Yang et al. (2023) for comparison. Based on these models, we design two settings: 1-shot reasoning and instruction tuning. In the first setting, we utilize the previously constructed example as ICL. In the second setting, we use parameter-efficient fine-tuning with LoRA Hu et al. (2021). The instruction data is generated by transforming the input and output from the training data.
In the closed-source LLMs, we use ChatGPT3.5-turbo [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo and GPT4o-mini Achiam et al. (2023) for comparison. We are unable to include this for all MSDiagnosis benchmarks due to the extraordinarily high cost of o1-preview inference. Based on these models, we employ two settings: 1-shot and prompting methods. In the first approach, we utilize the same example as previously used. In the second setting, we consider comparing method COT Wei et al. (2022) and method DAC Zhang et al. (2024). COT Wei et al. (2022) uses “Let’s think step by step” to enhance the model’s reasoning ability for task-solving. DAC Zhang et al. (2024) adopts a simple divide-and-conquer prompting strategy, where the input sequence is simply divided into multiple sub-inputs, which can enhance the reasoning performance of the LLM.
In other method, we mainly consider manual answering methods. Specifically, we randomly select 100 samples from MSDiagnosis and invite one college student (different from the annotation team in Section 3.2.2) to answer the questions. Our evaluation method for human results is consistent with other baselines, utilizing the automated assessment approach described in Section B.